The star of a distant solar system explodes as a supernova. At the moment of the explosion, an resting exploration spaceship is 15 AU away from the shock wave. The shock wave of the explosion travels with 25000 km/s towards the spaceship. To save the crew, the spacecra makes use of a special booster that uniformly accelerates at 150 m/s2 in the opposite direction. Determine if the crew manages to escape from the shock wave. (Neglect relativistic eects.)


1
Expert's answer
2021-03-23T16:57:40-0400
"t=\\frac{v}{a}=166667\\ s\\\\d=v _\ns\n\u200b\t\n t=4.167\u22c510^{12}m=27.85\\ AU"

So, only "27.85-15=12.85 AU" left for the spaceship to develop the speed of 25000 km/s. Is it possible? Find the distance required for the ship to accelerate from rest to 25 000 000 m/s with acceleration of 150 m/s2:

"d'=\frac{v^2}{2a}=2.08\cdot10^{12}m=13.9\ AU"

Therefore, when the shockwave will travel 27.85 AU from the star at 25 000 km/s, the speed of the spaceship will be lower than 25 000 km/s. They cannot escape.